Could you please answer the following. "A rectangle has a perimeter of 38 feet and an area of 48 square feet. What are the dimensions of the rectangle"?
Answer by oberobic(2304)

2(L+W) = 38
L*W = 48
so
W = 48/L
.
substitute
.
2(L+48/L) = 38
2L +96/L = 38
.
multiply both sides by L to eliminate the fraction
.
2L^2 +96 = 38L
.
2L^2 -38L +96 = 0
.
divide by 2
.
L^2 -19L +48 = 0
.
factor
.
(L-3)(L-16) = 0
.
L = 3 or L=16
so
W = 16 or 3, respectively
.
Done.